If you have a box of 10 cm of edge, how many grains of sand would fit in this box?


1
Expert's answer
2021-02-23T10:03:23-0500

A grain of sand weighs about


"m=10^{-6}g"

"N=\frac{V}{V_g}=\frac{V\rho}{m}\\N=\frac{10\cdot1.45}{10^{-6}}=1.45\cdot10^{7}"